We are continuing to monitor for any further issues.
Posted Apr 26, 2025 - 20:38 BST
Monitoring
All services are back online. Damage was found on a local fibre and has been fixed. We will continue to monitor the services.
Posted Apr 26, 2025 - 14:09 BST
Update
Our engineers continue to work on site to bring the services live again. More updates to follow.
Posted Apr 26, 2025 - 12:58 BST
Update
Engineer ETA on site will be 12:30pm. More updates to follow when they are on site.
Posted Apr 26, 2025 - 11:10 BST
Update
Our field engineers are heading to the area now, ETA to be confirmed
Posted Apr 26, 2025 - 11:06 BST
Investigating
MSO Ticket Reference : We are currently experiencing an incident impacting a small number of customer in the Mayfair area of London. Preliminary investigations are underway with our engineers to determine the root cause of this network incident. At this stage we are unable to issue an ERT until our engineers have completed further diagnostics.